If Peter's age plus Paul's age = 34
and Peter's age plus Mary's age =42
and Paul's age plus Mary's age =24
How old are Peter, Paul and Mary

Answer by TimothyLamb(4379)

x = age of Peter
y = age of Paul
z = age of Mary
---
x + y = 34
x + z = 42
y + z = 24
---
put the system of linear equations into standard form
---
2x + 2y = 70
x - 2y = 8
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 26
y = 8
z = 16
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---